The isotope 143Pr undergoes beta decay with a half-life of 13.57 days. What fraction of 143Pr has decayed after 11.6 days?
Answer #1

Radio active edecay is always first order reaction.

Lambda = 0.693 / 13.57days = 0.0512 days-1

Lambda = (2.303 / t ) Log (a / a-x)

0.0512 = (2.303 / 11.6 ) Log(1/1-x)

1/(1-x) = 1.81

1-x = 0.552

x = 0.448

So fraction of decay after 11.6 hours = 44.8 %